Fermi National Laboratory


Accelerator Update for September 26, 2005 - September 28, 2005

Monday September 26
The day shift began with Operations monitoring store 4401 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, and with the MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Operations began preparing the Recycler at 9:28 AM to stash antiprotons. The stash ended and Pbar resumed stacking at 10:36 AM.

Pbar experts began a reverse proton study at 11:02 AM.

Operations established beam to MCenter at 11:31 AM

The Pbar study ended at 4 PM. Pbar resumed stacking.

At 4:28 PM, two Main Injector (MI) RF stations tripped off (MIRF13 & 14). They reset okay, but then at 4:31 PM MIRF14 tripped off for the first of four times in nine minutes. It reset each time.

Operations terminated store 4401 at 6:01 PM.

Operations put the Tevatron (TeV) into shot setup at 6:15 PM.

Operations established the Recycler only store 4402 at 8:30 PM with an initial luminosity of 119.35E30. Pbar and MCenter resumed normal operations.

The Recycler began preparing to stash at 10:59 PM.

Tuesday September 27
The midnight shift began with Operations monitoring store 4402, with Pbar and the Recycler stashing, and with MiniBooNE and NuMI taking beam.

The Recycler stash ended and MCenter resumed taking beam around 12:08 AM. Pbar didn't resume stacking until 12:27 AM.

The Recycler began preparing to stash at 4:18 AM. The Recycler stash ended and Pbar and MCenter resumed normal operations at 5 AM.

The Recycler began preparing to stash at 8:49 AM. The Recycler stash ended at 10:10 AM.

An operator began preparing the Pbar tunnels at 10:14 AM for an access to repair a Debuncher RF station (DRF1-3).

MCenter resumed taking beam at 10:22 AM.

The Pbar access ended and Pbar resumed stacking at 12:28 PM.

At 12:54 PM, Operations reported that Pbar had lost its stack. Operators investigated and discovered a problem with an Accumulator quadrupole power supply (A:LQ). Experts replaced a voltage regulator and had the supply back up by 2:42 PM.

Pbar resumed stacking 3:13 PM.

At 8:14 PM, the TeV suffered a sector D4 quench. Operations investigated.

Operators prepared the CDF and D0 collision halls for access at 9:35 PM.

At 10:22 PM, Operations reported that a TeV sector A1 vacuum problem had cascaded into sector F0 and into the P1 transfer line. This put a halt to stacking.

Accelerator Overview

CDF came out of access at 10:23 PM. D0 never went in on access.

Operations reported at 10:43 PM that they had found a TeV corrector magnet power supply (HC48) that had a fuse fault and a 40 amp deviation. This occurred coincidentally with the quench. Operators replaced its dipole function generator (DFG).

Operations reported at 11:23 PM that the P1 vacuum had improved. By 11:38 PM it was good enough for Pbar to resume stacking.

Wednesday September 28
The midnight shift began with Operations monitoring the stack, with the TeV recovering from a quench, and with MiniBooNE and NuMI taking beam.

At 12:30 AM, Operations reported more regulation problems with HC48. An expert advised Operations to replace the DFG again.

The Recycler began stashing at 12:46 AM. The stash ended and Pbar resumed stacking at 1:35 AM. The Recycler set a New Record for Stored Antiprotons with a level of 247E10.

Operations put the TeV into dry squeeze at 1:17 AM.

Operations prepared the TeV at 1:47 AM for a proton only store.

Operations put the TeV through a wet squeeze at 3:02 AM.

Operations put the TeV into shot setup at 3:39 AM.

At 5:38 AM, after all the protons and antiprotons were loaded and being accelerated, the TeV suffered a sector F4 quench. Later a TeV expert reported that the cause of the quench was due to the new orbit stabilizer program; the new beta star settings were never copied to the operational table.

Operations reported at 5:48 AM that nearly all the vacuum pumps around sector A0 tripped off after the quench.

The Plans for Wednesday
The plans for today are to go into a Pbar only shot setup around noon. The Run Coordinator will allow the TeV experts to conduct studies until noon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/28/2005   email Fermilab